# Measurement of $\Sigma^{+}$ transverse polarization in $e^{+}e^{-}$ collisions at $\sqrt{s}=3.68-3.71$ GeV

### v2

@@ 6 Summary
In summary, based on $e^{+}e^{-}$ collision data corresponding to an integrated luminosity of 652.1 pb-1 collected with the BESIII detector at the BEPCII collider, we present an energy-dependent measurement of transverse polarization, relative phase and the modulus ratio of the form factors of $\Sigma^{+}$ hyperon in the $e^{+}e^{-}\to\Sigma^{+}\bar{\Sigma}^{-}$ reaction. For the first time, the phase of the $\Sigma^{+}$ hyperon electromagnetic form factors is explored in a higher range of four-momentum transfer above $q^{2}>13.5~{}\rm{GeV}^{2}$. No polarization is evident at each energy point. The comparison between this work and previous measurements BESIII:2023sgt ; BESIII:2023ynq is shown in figure 5. $\Delta\Phi$ is less than zero for $J/\psi$ decay and $\sqrt{s}=3.71$ GeV, and greater than zero at other energy points, which implies that there may be at least one $\Delta\Phi=0$ rad between these energy points. Such an evolution is important input for understanding its asymptotic behavior and the dynamics of baryons similar to the $\Lambda$ hyperon Mangoni:2021qmd . These results are important to understand the production mechanism of the $\Sigma^{+}$-$\bar{\Sigma}^{-}$ pairs at different energy points.

### v3

@@ 6 Summary
In summary, based on $e^{+}e^{-}$ collision data corresponding to an integrated luminosity of 652.1 pb-1 collected with the BESIII detector at the BEPCII collider, we present an energy-dependent measurement of transverse polarization, relative phase and the modulus ratio of the form factors of $\Sigma^{+}$ hyperon in the $e^{+}e^{-}\to\Sigma^{+}\bar{\Sigma}^{-}$ reaction. For the first time, the phase of the $\Sigma^{+}$ hyperon electromagnetic form factors is explored in a higher range of four-momentum transfer above $q^{2}>13.5~{}\rm{GeV}^{2}$. No polarization is evident at each energy point. These results are important to understand the production mechanism of the $\Sigma^{+}$-$\bar{\Sigma}^{-}$ pairs at different energy points.